OpenStudy (anonymous):

Okay, I have a test in 8 Hours and I don't know how to do this. Can someone explain the steps for me to memorize. 1. A company wishes to manufacture a box with a volume of 36 ft3 that is open on top and is twice as long as it is wide. (a) Let x be the length of the box (so the width is 2x), and let y be the height of the box. First write the surface area of the box as a function of x and y. Then use the fact that the volume is 36 cubic feet to write the surface area as a function of x only. (b) Find the dimensions of the box that will lead to the minimum surface area. (c) What is the minimum surface area?

OpenStudy (anonymous):

Here is the solultion: a.) Surface Area = 2x^2 + 6xy = 2x^2 = 108/x b.) x=3 y=2 c.) 30 square feet. I need to learn the steps to solve these

OpenStudy (anonymous):

my book has it laid out in a bunch of sections and its gonna take too long to cipher through it all.

hartnn (hartnn):

vol=36=(2*x)*x*y=>y=18/(x*x) surface area=(2*x)*x+2*y*x+2*y*(2*x)=2*x*x+6*x*y now put y in this......then differentiate w.r.t x and equate it to zero for maximization...to get x=3*cuberoot(2)....once u know x,u know everything.....

hartnn (hartnn):

sorry minimization...

hartnn (hartnn):

sorry to get x=3

OpenStudy (anonymous):

oh sorry this is calc, cant do anything

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

:(

ganeshie8 (ganeshie8):

did u finish with part a ?

ganeshie8 (ganeshie8):

@MercerC

OpenStudy (anonymous):

What does he mean by differentiate w.r.t. x?

ganeshie8 (ganeshie8):

it just means taking derivative

OpenStudy (anonymous):

Followed it up until that part

ganeshie8 (ganeshie8):

ok. so from part a, you have the surface area function in x & y SA = 2x^2 + 6xy ---------------------------(1) surface area in x : SA = 2x^2 + 108/x -------------------------(2)

ganeshie8 (ganeshie8):

you want to minimize (2), right ?

hartnn (hartnn):

yes,taking derivative....and you get 4*x-108/(x*x)...when u equate this to 0,u get x=3.

hartnn (hartnn):

now is there any part u did not get?

OpenStudy (anonymous):

Sorry, I'm not as quick. I'm at Put y into this....which I'm guessing you mean the surface area equation which is for x

OpenStudy (anonymous):

2*x*x + 6 * (18/x*x) * x ?

hartnn (hartnn):

yes,this gives 2x*x+108/x as u expected(given in ans)

OpenStudy (anonymous):

so I'm on a financial calculator who do I get the cube root of something. the 108/x is throwing me off

OpenStudy (anonymous):

how*

OpenStudy (anonymous):

18/x^2....what do I need to plug in or how do I need to get rid of the x^2 here?

hartnn (hartnn):

are u struck?if yes,where?

OpenStudy (anonymous):

I don't understand the 108/x

OpenStudy (anonymous):

I get that its easy math for you guys, but I'm plugging stuff in on this calculator and jsut can't actually come up with it, which means I'm not gonna get it on the test.

OpenStudy (anonymous):

Then I know how to diffrentiate for fxx(x,y) and fyy(x,y) to get 3 and 2 if those are the steps

OpenStudy (anonymous):

I'm stupid....6*18 = 108

OpenStudy (anonymous):

A Mathematica part solution with comments is attached.

OpenStudy (anonymous):

hartnn, so when I differentiate. I use 2x^2 + 6xy?

OpenStudy (anonymous):

for x that is

OpenStudy (anonymous):

for part b

OpenStudy (anonymous):

....getiing it i think

OpenStudy (anonymous):

nope i'm not. please help with part b

OpenStudy (anonymous):

Finally followed your steps. thanks. It took me a few times to find what I needed, but I got it.

hartnn (hartnn):

finally!!